OpenStudy (anonymous):

Find the slope-intercept form of the equation of a line with slope 1 containing point (-4, 10). Question 10 options: y = x + 6 y = x + 14 y = -x + 14 y = 14x + 1

OpenStudy (hoblos):

y-10 = x +4 y = x+14